OpenStudy (anonymous):

A cylinder is full at 385 cm3 and has a radius of 4.5 cm. It currently contains 225 cm3 of water. What will the difference in the height of the water be when the cylinder is full? Use 3.14 to approximate pi, and round your answer to one decimal place. _______cm

OpenStudy (anonymous):

\[Area = 2 \times \pi \times r^2 \times h\]

OpenStudy (anonymous):

where 'r' is 4.5cm pi is 3.14

OpenStudy (anonymous):

OOPS! ITS VOLUME NOT AREA SORRY!

OpenStudy (anonymous):

\[V = 2 \times \pi \times r^2 \times h\]

OpenStudy (anonymous):

its okay hun i already took the test and got 80% so ya lol

OpenStudy (anonymous):

You got this sum correct?

OpenStudy (anonymous):

Volume 1 = 385 Volume 2 = 225

OpenStudy (anonymous):

Ya :)

OpenStudy (anonymous):

So what was the answer??